The gravitational attraction between a 20 kg cannonball and a marble separated center-to-center by 30 cm is 1.48 x 10 ^ -10 N. Compute the mass of the marble.


1
Expert's answer
2021-01-10T18:09:34-0500

The gravitational attraction between two masses "m_1,m_2" is given as follows:


"F = G\dfrac{m_1m_2}{r^2}"

where "m_1 = 20kg", "m_2" is the mass of the marble, "r = 30cm = 0.3m" is the distance between centers, and "G = 6.67\times 10^{-11}m^3/(kg\cdot s^2)" . Expressing "m_2" and substituting "F = 1.48\times 10^{-10}N", obtain:


"m_2 = \dfrac{r^2F}{m_1G}\\
m_2 = \dfrac{0.3^2\cdot 1.48\times 10^{-10}}{30\cdot 6.67\times 10^{-11}} \approx 6.66\times 10^{-3}kg"

Answer. "6.66\times 10^{-3}kg".